OpenStudy (anonymous):

What is the value of the coefficient "a" when the quadratic equation y = (3x - 4)(2x - 1) is written in standard form?

OpenStudy (anonymous):

6

OpenStudy (mertsj):

Do you know what standard form is?

OpenStudy (anonymous):

foil that bad mofo!

OpenStudy (anonymous):

iknow what it is buti forgot how to do it

OpenStudy (anonymous):

\( (3 x - 4) (2 x - 1) = 4 - 11 x + 6 x^2 \) comparing with \(ax^2 +bx+c \Rightarrow a = 6 \)

OpenStudy (anonymous):

Man i was right and didn't get any medals 8(